OpenStudy (anonymous):

-3(2x-4) can I simplify the -3 & the -4 regardless that the -4 is in the parenthesis

OpenStudy (anonymous):

yes you may re arrange them, mybe like this -3(2x-4) =3(-2x+4) or =3(4-2x) is that maybe what you mean

OpenStudy (anonymous):

Yes thank you. The parenthesis threw me off

OpenStudy (anonymous):

So I combined the like terms -3 -4+2x which I got 7+2x

OpenStudy (ajprincess):

You can not combine like that. First u have to remove the parentheses using distributive property. That is a(b+c)=a*b+a*c Then only you have to combine the like terms.

Directrix (directrix):

@ritamo24 Your Question: -3(2x-4) can I simplify the -3 & the -4 regardless that the -4 is in the parenthesis Answer: No. The -4 you see within the parenthesis is not really a -4. It is a (-4) multiplied by a (-3).